# Long-Tailed Classification Based on Coarse-Grained Leading Forest and Multi-center Loss

### v2

@@ Abstract
Long-tailed (LT) classification is an unavoidable and challenging problem in the real world. Most existing long-tailed classification methods focus only on solving the class-wise imbalance while ignoring the attribute-wise imbalance. The deviation of a classification model is caused by both class-wise and attribute-wise imbalance. Due to the fact that attributes are implicit in most datasets and the combination of attributes is complex, attribute-wise imbalance is more difficult to handle. For this purpose, we proposed a novel long-tailed classification framework, aiming to build a multi-granularity classification model by means of invariant feature learning. This method first unsupervisedly constructs Coarse-Grained forest (CLF) to better characterize the distribution of attributes within a class. Depending on the distribution of attributes, one can customize suitable sampling strategies to construct different imbalanced datasets. We then introduce multi-center loss (MCL) that aims to gradually eliminate confusing attributes during feature learning process. The proposed framework does not necessarily couple to a specific LT classification model structure and can be integrated with any existing LT method as an independent component. Extensive experiments show that our approach achieves state-of-the-art performance on both existing benchmarks ImageNet-GLT and MSCOCO-GLT and can improve the performance of existing LT methods. Our codes are available on GitHub: https://github.com/jinyery/cognisance
@@
Imbalanced learning, long-tailed learning, Coarse-Grained forest, invariant feature learning, multi-center loss.
@@ I Introduction
Current research mostly focuses on solving the class-wise long-tailed problem, where resampling [10, 11] or loss reweighting [12, 13] methods are often used to rebalance the training process. However, most of these methods sacrifice the performance of the head class to improve the performance of the tail class, which is like playing a performance seesaw, making it difficult to fundamentally improve the performance for all classes. Apart from resampling and reweighting, there are some methods (e.g., [14, 15]) believe that data imbalance does not affect feature learning, so the training of the model is divided into two stages of feature learning and classifier learning. However, this adjustment is only a trade-off between accuracy and precision [9, 16], and the confusion areas of similar attributes in the feature space learned by the model do not change. As depicted in Fig 2, it is the attribute-wise imbalance that leads to spurious correlations between the head attributes and a particular category. This means that these head attributes correspond to the spurious features (i.e., a confusing region in the feature space) of the category.
@@
This paper proposes a framework termed as Cognisance, which is founded on Coarse-grained Leading Forest and Multi-center Loss. Cognisance handles class-wise and attribute-wise imbalance problems simultaneously by constructing different environments 111 Environments are the data subsets sampled from existing image datasets to reflect both class-wise and attribute-wise imbalance. , in which we design a new sampling method based on a structure of samples called Coarse-Grained Leading Forest (CLF). A CLF is constructed by unsupervised learning that characterizes the attribute distributions within a class and guides the data sampling in different environments during training process. In the experimental setup of this paper, two environments are constructed, one of which is the original environment without special treatment, and the distributions of categories and attributes in the other are balanced. In order to gradually eliminate confusing pseudo-features during the training process, we design a new metric learning loss, Multi-Center Loss (MCL), which is inspired by [9] and [19], extends the center loss to its Invariant Risk Minimization (IRM) version. MCL enables our model to better learn invariant features and further improves the robustness when compared to its counterparts in [9] and [19]. In addition, the proposed framework is not coupled with a specific backbone model or loss function, and can be seamlessly integrated into other LT methods for performance enhancement.
@@ II-A Long-Tailed Classification
The key challenge of long-tailed classification is to effectively deal with the imbalance of data distribution to ensure that excellent classification performance can be achieved both between the head and the tail. Current treatments for long-tailed classification can be broadly categorized into three groups [1]: 1) Class Re-balancing, which is the mainstream paradigm in long-tailed learning, aims to enhance the influence of tail samples on the model by means of re-sampling [20, 11, 21, 22], re-weighting [23, 24, 25, 26, 27] or logit adjustment [28] during the model training process, and some of these methods [14, 15] consider that the unbalanced samples do not affect the learning of the features, and thus divide feature learning and classifier learning into two phases, and perform operations such as resampling only in the classifier learning phase. 2) Information Augmentation based approaches seek to introduce additional information in model training in order to improve model performance in long-tailed learning. There are two approaches in this method type: migration learning [29, 30, 31] and data augmentation [32, 33, 34]. 3) Module Improvement seeks to improve network modules in long-tailed learning such as RIDE [35] and TADE [36], both of which deal with long-tail recognition independent of test distribution by introducing integrated learning of multi-expert models in the network.
@@
Remark: In practical applications, the formula does not impose the untangling assumption, i.e., it does not assume that a perfect feature vector $z=[z_{c};z_{a}]$ can be obtained, where $z_{a}$ and $z_{c}$ are separated.
@@ III-A Coarse-Grained Leading Forest
1) Compute the distance matrix between the sample points in the dataset X using an arbitrary distance metric. Then compute the density of sample point $i$ is computed as in Eq. (3):
@@
where $\boldsymbol{I}=\left\{1,2,...,N\right\}$ is the index set of dataset $\boldsymbol{X}$, $d_{rd}$ is the radius of the hypersphere for computing density centered at $\boldsymbol{x}_{i}$, $\boldsymbol{O}_{i}$ is the set of samples lie outside of the hypersphere that do not contribute to $\rho_{i}$.
2) Sort the samples in descending order according to the density value, denoted as $\boldsymbol{S}$ the indices of the sorted data, i.e., $\boldsymbol{S}_{i}$ is the index of the data point with the $i$-th largest density value.
3) Form Coarse-Grained node on the points in $\boldsymbol{S}$ sequentially, which relies on the hyperparameter $d_{rn}$. If the data point $\boldsymbol{S}_{i}$ has not yet been merged into a Coarse-Grained node, then the data point $\boldsymbol{S}_{i}$ and the points within $d_{rn}$ distance from it are formed into a Coarse-Grained node, i.e.:
@@ 14       $l_{i}$ = find_leader($i$, $d_{rn}$);
4) Find the leading node of the newly constructed Coarse-Grained node. Here, the problem is transformed into finding the leading node of the prototype $\boldsymbol{S}_{i}$ of a Coarse-Grained node, denoted as $l_{i}$, and then $\boldsymbol{C}(l_{i})$ is assigned as the leading node of $\boldsymbol{C}(i)$. The process of finding the leading node $\boldsymbol{C}(l_{i})$ of $\boldsymbol{C}(i)$ can be formulated as Eq. (5):
where $\boldsymbol{O}_{S_{i}}$ is the set of nodes whose distance from $\boldsymbol{S}_{i}$ exceeds $d_{rd}$. Note that $l_{i}$ may not exist, and when $l_{i}$ is not found, the Coarse-Grained node $\boldsymbol{C}(i)$ is the root of a leading tree in the whole Coarse-Grained leading forest. Also, since the density of nodes in $\boldsymbol{S}$ is decreasing, when $l_{i}$ is found, it must have been processed and already merged into $\boldsymbol{C}(i)$ since $l_{i}$ is denser.
@@ III-B Sampling based on CLF
After the samples of each attribute is determined, we can refer to the idea of class-wise resampling [14] to resample on the attributes (compute the probability of each sampled objects from class $j$ ), as shown in (6).
@@
where $j$ represents the class index, $n_{j}$ is the number of samples of class $j$, and $C$ the total number of classes. $q\in[0,1]$ is a user-defined parameter, and class-wise balanced sampling is performed when $q=0$.
@@
1 weights = initial_vector(default=0, size=||||clf|normal-|||);
@@ III-C Multi-Center Loss
$\begin{split}&\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i\in e}L_{cls}\big{(}f(x_{i}^{e};\ \theta),\ y_{i}^{e};\ w\big{)}\\ \mathrm{s.t.}&\quad\theta\in\mathop{\arg\min}\limits_{\Theta}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i\in e}{\|f(x_{i}^{e};\ \Theta)-\mathcal{C}(x_{i})\|}_{2},\end{split}$ (7)
@@
$\begin{split}\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i\in e}L_{mc}&=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i\in e}L_{cls}+\alpha\cdot L_{IFL}\\ &=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i\in e}L_{cls}\big{(}f(x_{i}^{e};\ \theta),\ y_{i}^{e};\ w\big{)}\\ &\quad\ +\alpha\cdot{\|f(x_{i}^{e};\ \theta)-\mathcal{C}(x_{i})\|}_{2}.\end{split}$ (8)

### v3

@@ Abstract
Long-tailed (LT) classification is an unavoidable and challenging problem in the real world. Most existing long-tailed classification methods focus only on solving the class-wise imbalance while ignoring the attribute-wise imbalance. The deviation of a classification model is caused by both class-wise and attribute-wise imbalance. Due to the fact that attributes are implicit in most datasets and the combination of attributes is complex, attribute-wise imbalance is more difficult to handle. For this purpose, we proposed a novel long-tailed classification framework, aiming to build a multi-granularity classification model by means of invariant feature learning. This method first unsupervisedly constructs Coarse-Grained Leading Forest (CLF) to better characterize the distribution of attributes within a class. Depending on the distribution of attributes, one can customize suitable sampling strategies to construct different imbalanced datasets. We then introduce multi-center loss (MCL) that aims to gradually eliminate confusing attributes during feature learning process. The proposed framework does not necessarily couple to a specific LT classification model structure and can be integrated with any existing LT method as an independent component. Extensive experiments show that our approach achieves state-of-the-art performance on both existing benchmarks ImageNet-GLT and MSCOCO-GLT and can improve the performance of existing LT methods. Our codes are available on GitHub: https://github.com/jinyery/cognisance111©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Imbalanced learning, long-tailed learning, coarse-grained leading forest, invariant feature learning, multi-center loss.
@@ I Introduction
Current research mostly focuses on solving the class-wise long-tailed problem, where resampling [10] [11], loss reweighting [12] [13], tail class augmentation [14], and Gaussian perturbed feature [15] [16] are used to rebalance the training process. However, most of these methods sacrifice the performance of the head class to improve the performance of the tail class, which is like playing a performance seesaw, making it difficult to fundamentally improve the performance for all classes. Apart from resampling and reweighting, there are some methods (e.g., [17, 18]) believe that data imbalance does not affect feature learning, so the training of the model is divided into two stages of feature learning and classifier learning. However, this adjustment is only a trade-off between accuracy and precision [9, 19], and the confusion areas of similar attributes in the feature space learned by the model do not change. As depicted in Fig 2, it is the attribute-wise imbalance that leads to spurious correlations between the head attributes and a particular category. This means that these head attributes correspond to the spurious features (i.e., a confusing region in the feature space) of the category.
@@
This paper proposes a framework termed as Cognisance, which is founded on Coarse-grained Leading Forest and Multi-center Loss. Cognisance handles class-wise and attribute-wise imbalance problems simultaneously by constructing different environments 222 Environments are the data subsets sampled from existing image datasets to reflect both class-wise and attribute-wise imbalance. , in which we design a new sampling method based on a structure of samples called Coarse-Grained Leading Forest (CLF). A CLF is constructed by unsupervised learning that characterizes the attribute distributions within a class and guides the data sampling in different environments during training process. In the experimental setup of this paper, two environments are constructed, one of which is the original environment without special treatment, and the distributions of categories and attributes in the other are balanced. In order to gradually eliminate confusing pseudo-features during the training process, we design a new metric learning loss, Multi-Center Loss (MCL), which is inspired by [9] and [22], extends the center loss to its Invariant Risk Minimization (IRM) version. MCL enables our model to better learn invariant features and further improves the robustness when compared to its counterparts in [9] and [22]. In addition, the proposed framework is not coupled with a specific backbone model or loss function, and can be seamlessly integrated into other LT methods for performance enhancement.
@@
  • We propose a novel framework to deal with noise in long-tailed datasets, in which we design a new loss function and a noise selection strategy based on CLF. The framework is validated on two long-tailed noise datasets, and achieves encouraging performance.
@@ II-A Long-Tailed Classification
The key challenge of long-tailed classification is to effectively deal with the imbalance of data distribution to ensure that excellent classification performance can be achieved both between the head and the tail. Current treatments for long-tailed classification can be broadly categorized into three groups [1]: 1) Class Re-balancing, which is the mainstream paradigm in long-tailed learning, aims to enhance the influence of tail samples on the model by means of re-sampling [23, 11, 24, 25], re-weighting [26, 27, 28, 29, 30] or logit adjustment [31, 15] during the model training process, and some of these methods [17, 18] consider that the unbalanced samples do not affect the learning of the features, and thus divide feature learning and classifier learning into two phases, and perform operations such as resampling only in the classifier learning phase. 2) Information Augmentation based approaches seek to introduce additional information in model training in order to improve model performance in long-tailed learning. There are two approaches in this method type: migration learning [32, 33, 34] and data augmentation [35, 36, 37]. 3) Module Improvement seeks to improve network modules in long-tailed learning such as RIDE [38] and TADE [39], both of which deal with long-tailed recognition independent of test distribution by introducing integrated learning of multi-expert models in the network. In addition, SHIKE [40] propose depth-wise knowledge fusion to fuse features between different shallow parts and the deep part in one network for each expert, which makes experts more diverse on representation.
@@
Remark 1: In practical applications, the formula does not impose the untangling assumption, i.e., it does not assume that a perfect feature vector $z=[z_{c};z_{a}]$ can be obtained, where $z_{a}$ and $z_{c}$ are separated.
@@ III-A Coarse-Grained Leading Forest
1) Calculate the distance matrix and sample density: First calculate the distance matrix between the samples in the dataset $\boldsymbol{X}$ using an arbitrary distance metric, and then calculate the density of each sample point $i$ with reference to Eq. (3).
@@
[…]
2) Creating coarse-grained nodes: First arrange the samples in order of decreasing density values, denoted as $\boldsymbol{S}$ the indices of the sorted data, i.e., $\boldsymbol{S}_{i}$ is the index of the data point with the $i$-th largest density value. Then iterate through the samples in $\boldsymbol{S}$ sequentially, and if the data point $\boldsymbol{S}_{i}$ has not yet been merged into a coarse-grained node, then the data point $\boldsymbol{S}_{i}$ and the points within $d_{rn}$ distance from it are formed into a coarse-grained node, i.e.:
@@
3) Finding the leading node: Whenever a new coarse-grained node is created, a lead node is found for it. Here, the problem is transformed into finding the leading node of the prototype $\boldsymbol{S}_{i}$ of a coarse-grained node, denoted as $l_{i}$, and then $\boldsymbol{C}(l_{i})$ is assigned as the leading node of $\boldsymbol{C}(i)$. The process of finding the leading node $\boldsymbol{C}(l_{i})$ of $\boldsymbol{C}(i)$ can be formulated as Eq. (5):
where $\boldsymbol{O}_{S_{i}}$ is the set of nodes whose distance from $\boldsymbol{S}_{i}$ exceeds $d_{rd}$. Note that $l_{i}$ may not exist, and when $l_{i}$ is not found, the coarse-grained node $\boldsymbol{C}(i)$ is the root of a leading tree in the whole coarse-grained leading forest. Also, since the density of nodes in $\boldsymbol{S}$ is decreasing, when $l_{i}$ is found, it must have been processed and already merged into $\boldsymbol{C}(i)$ since the density of $l_{i}$ is higher.
Time Complexity. The time complexity of calculating the distance matrix between sample points is $O(N^{2})$, where $N$ is the number of sample points. Next, calculating the density for each sample point also has a complexity of $O(N^{2})$. Following this, the time complexity for sorting the densities in descending order is $O(N\log N)$. Within the main loop, each iteration may need to check all unvisited points to determine which points are within the distance $d_{rn}$, with the worst-case complexity being $O(N)$. Similarly, when finding the node with the highest density within distance $d_{rd}$, the worst-case complexity is also $O(N)$. Overall, the primary time-consuming operations are the distance and density calculations, resulting in an overall time complexity of $O(N^{2})$.
Space Complexity. The main cost arises from storing the distance matrix between sample points, with a complexity of $O(N^{2})$. Additionally, storing the density information for each sample point, the sorting indices, and the boolean vector all have complexities of $O(N)$. Consequently, the overall space complexity of the algorithm is also $O(N^{2})$.
Although the time and space complexity of the algorithm is relatively high, it spends mainly on the computation and storage of the distance matrix, operations that can be optimized by distributed or parallel methods. A number of acceleration schemes for density-peak clustering have been proposed. For example, in another work of ours, FaithPDP [43], which is specifically optimized for the construction of the leading tree structure, a parallel solution is proposed, while instead of storing the entire distance matrix, only two vectors and a tall matrix need to be stored. These improvements significantly improve the efficiency of the algorithm, especially when dealing with large-scale datasets.
@@ III-B Sampling based on CLF
After the samples of each attribute is determined, we can refer to the idea of class-wise resampling [17] to resample on the attributes (compute the probability of each sampled objects from class $j$ ), as shown in Eq. (6).
@@
where $j$ represents the class index, $n_{j}$ is the number of samples of class $j$, and $C$ the total number of classes. And when performing intra-class sampling, $j$ represents the attribute index, $n_{j}$ is the number of samples of attribute $j$, and $C$ the total number of attributes. $q\in[0,1]$ is a user-defined parameter, and balanced sampling is performed when $q=0$.
Remark 2: Both inter-class and intra-class resampling refer to Eq. (6), except that the meaning of the symbols in the equations is different for the two types of sampling.
@@
Remark 3: The weights calculated for each sample in the above steps are only relative values (i.e., do not necessarily sum to one) and need to be normalized after all sample weights have been calculated.
@@
1 weights = initial_vector(default=0, size=||||clf||||);
@@ III-C Multi-Center Loss
$\begin{split}&\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{cls}\big{(}f(x_{i}^{e};\ \theta),\ y_{i}^{e};\ w\big{)}\\ \mathrm{s.t.}&\quad\theta\in\mathop{\arg\min}\limits_{\Theta}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}{\|f(x_{i}^{e};\ \Theta)-\mathcal{C}(x_{i})\|}_{2},\end{split}$ (7)
@@
$\begin{split}\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{mc}&=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{cls}+\alpha\cdot L_{IFL}\\ &=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{cls}\big{(}f(x_{i}^{e};\ \theta),\ y_{i}^{e};\ w\big{)}\\ &\quad\ +\alpha\cdot{\|f(x_{i}^{e};\ \theta)-\mathcal{C}(x_{i})\|}_{2}.\end{split}$ (8)
@@ 13       $\mathcal{E}=\text{EnvConstruct}(\{(q_{cls}^{e},q_{attr}^{e})\},clf)$;
III-E Noise Selection
Identifying noise in long-tailed datasets is a highly challenging task [44]. Due to the low frequency of tail samples in long-tailed datasets, accurately distinguishing genuine tail categories from noise becomes extremely difficult. In such cases, the number of noise samples may be comparable to or even exceed that of tail category samples, thus increasing the complexity of model training and application. Especially, when using resampling methods, the resampling process may lead to an increase in the proportion of noise samples in the dataset, thereby negatively impacting model training.
Figure 7: Distribution of noise samples in clustering under different feature dominances; noise samples are more peripheral under invariant features.
To tackle this challenge, we propose a noise selection algorithm based on the Cognisance. The algorithm has two main motivations: First, by learning invariant features from the data, normal and noise samples can be clearly distinguished in the feature space. As shown in Fig. 7, when clustering based on pseudo-features, the “black swan” appears to be very outlier due to its rarity. If outlier-ness is simply used as the standard for identifying noise samples, the “black swan” is likely to be misidentified as noise, while the “white rooster” might be misidentified as a normal sample. However, by removing the pseudo-feature “color” and instead using the invariant feature “shape” as the main basis for identification, the “black swan” and “white swan” will cluster together, while the “white rooster” will appear as an outlier due to the shape difference.
Next, based on visualizations from previous work, we observed the distribution characteristics of noise samples in guiding forests. As shown in the red boxes of Fig. 3 and Fig. 4, noise samples are often located at the bottom of leading trees and, due to their outlier nature, the number of nodes forming a tree is usually sparse. Based on these quantity and location characteristics, we designed the core noise selection strategy:
1) Cluster Size-Based Noise Labeling: For small clusters of samples, if the number of samples in the cluster is less than the threshold $N_{\text{min}}$, all samples in the cluster will be directly labeled as noise. This criterion is designed based on the characteristic that noise samples typically form smaller clusters in the coarse-grained leading tree.
2) Depth and Layer-Based Noise Selection: In addition, two parameters $N_{d}$ and $N_{l}$ are set, representing the depth at which noise selection starts in each coarse-grained leading tree and the number of layers selected from bottom to top, respectively. For example, when $N_{l}=1$, all samples in the last layer of the coarse-grained leading tree are marked as noise. The setting of $N_{d}$ is to reduce the possibility of mislabeling normal samples as noise, as the deeper the sample is in the tree, the higher its “outlierness”, and the more likely it is to be a noisy sample. This criterion is based on the characteristic that noise samples are often located at deeper levels in the coarse-grained leading tree.
3) Density Percentile-Based Noise Filtering: Finally, to reduce the risk of selecting normal samples by mistake, this scheme introduces a filtering mechanism based on density percentiles at the end of the process. Density to some extent represents outliers, and by screening samples with lower density, the possibility of mislabeling normal samples can be reduced. Setting the density percentile threshold $P_{d}$ can limit the maximum number of labeled noise samples. For example, when $P_{d}=10\%$, the number of selected noise samples will not exceed 10% of the total number of samples in that category, and these noise samples are the 10% with the lowest density value among all samples. This step enhances the controllability of noise selection, preventing too many samples from being marked as noise.
To label a sample as noise, it is sufficient to satisfy either of the first two criteria (cluster size-based or depth and layer-based) and simultaneously satisfy the third criterion (density percentile-based).
To better highlight the quantity and location characteristics of noise samples, this paper introduces a new metric learning loss, the Multi-Center Triplet Loss (MCTL). This loss function not only brings samples of the same class closer to the center, but also pushes samples of different classes further apart, as shown in Eq. (9):
$\begin{split}&\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{mc}=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{cls}+\alpha\cdot L_{IFL}\\ &=\mathop{\min}\limits_{\theta,\ w}\sum\limits_{e\in\mathcal{E}}\sum\limits_{i}L_{cls}(g(f(x_{i}^{e};\ \theta);w)\ ,y_{i}^{e})\\ &\quad+\alpha\cdot\mathop{\max}(0,{\|f(x_{i}^{e};\ \hat{\theta})-C_{p}(x_{i}^{e})\|}_{2}\\ &\quad-{\|f(x_{i}^{e};\ \hat{\theta})-C_{n}(x_{i}^{e})\|}_{2}),\end{split}$ (9)
where $L_{IFL}$ is the invariant feature learning constraint loss, and $\alpha$ is the balancing parameter. This loss function is an IRM version of the triplet loss, which further enhances the robustness of the algorithm compared to the multi-center loss. As shown in Fig. 8, it is precisely because MCTL widens the distance between samples of different categories that the position of noisy samples can be closer to the bottom.
Figure 8: The CLF constructed based on the learned representations using different loss functions. The left figure shows results using ordinary cross entropy loss, while the right figure depicts results from MCTL. The red and blue boxes represent noise from “Jaguar” and “Lynx”, respectively. In the right, the noise samples are concentrated at the bottom.
We refer to this algorithm designed specifically for long-tailed noise datasets as the Cognisance+ framework. The specific process is shown in algorithm 4. Compared with the Cognisance framework, this method can further improve model performance on noisy datasets.
Output: backbone $f(\cdot;\theta)$, classifier $g(\cdot;w)$
1 Initialize: backbone $f(\cdot;\theta)$, classifier $g(\cdot;w)$;
2 for $M$ warm-up epochs do
       // Optimize the model for any loss
4
5 end for
6noises = initial_list();
7 $\{F_{y}\}=\textit{construct\_clf}(\{(x,y)\},\theta,w)$;
8 $\{e_{n}\}=\textit{construct\_env}(\{(q_{cls}^{e},q_{attr}^{e})\},\{F_{y}\},\textit{noises})$;
9 for $N$ epochs do
10       $\{(x^{e},y^{e})\}=\textit{data\_load}(\{e_{n}\})$;
11       $\{(C_{p}^{x},C_{n}^{x})\}=\textit{read\_centers}(\{F_{y}\},\{(x^{e},y^{e})\})$;
12       $\theta,w\leftarrow\hat{\theta},\hat{w}\in\text{argmin}_{\theta,w}L_{cls}(g(f(x;\theta);w),y)+\alpha\cdot L_{IFL}(f(x;\theta),\{(C_{p}^{x},C_{n}^{x})\})$;
13       $\{F_{y}\}=\textit{construct\_clf}(\{(x,y)\},\theta)$;
       // Iteratively update noise samples and environment.
[…]
15       $\{e_{n}\}=\textit{construct\_env}(\{(q_{cls}^{e},q_{attr}^{e})\},\{F_{y}\},\textit{noises})$;
16
17 end for
return $f(\cdot;\theta)$, $g(\cdot;w)$
Algorithm 4 Overall Framework Training Process